The people on coral Island by only juicing cloth. The CPO basket contains a quantities bought in 2013. The average household spent $34 on juice and $40 on cloth and 2013 when the price of juice was two dollars a bottle and the price of cloth was four dollars a yard. in 2014 juices four dollars a bottle and cloth is seven dollars a yard. Calculate the CPI basket and the percentage of the households for budget spent on juice in 2013


1
Expert's answer
2022-02-10T08:24:46-0500

Consumer Price Index

"CPI=\frac{C_t}{C_o}\times 100"

"C_t=current\ price"

"C_o=base\ price"

2013 is the base year in this case and the current year at the same time.

"\therefore C_t=(34+40)=74"

"C_o=74\ Since\ 2013\ is\ base\ and\ current\ year"

"\bold{CPI=\frac{74}{74}\times 100=1}"


Percentage spent on Juice

"Total\ budget=34+40=74"

"Expenditure\ on \ Juice=34"

"\% expenditure =\frac{34}{74}\times100=\bold{46\%}"